Find an equation for the line perpendicular to the line −8x − 8y = −8 having the same y-intercept as x − 2y = −14
Answer by stanbon(75887)

Find an equation for the line perpendicular to the line −8x − 8y = −8 having the same y-intercept as x − 2y = −14
-----------------------------------
Find the slope of x + y = 1
y = -x + 1
slope = -1
A perpendicular line will have slope = +1.
--------
Find the y-int of y = (1/2)x +7
y-int is (0,7)
------
Equation::
y = x + 7
---------------
Cheers,
Stan H.
------------